OpenStudy (anonymous):

WHICH ONE IS IT What value of 'b' would make 10y^2 + by + 3 factorable? a) 20 b) 23 c) 11 d) 12

jimthompson5910 (jim_thompson5910):

D = b^2 - 4ac ... start with the discriminant formula D = b^2 - 4*10*3 ... plug in a = 10, c = 3 D = b^2 - 120 ... multiply

jimthompson5910 (jim_thompson5910):

Now plug in each answer choice and see if you get a perfect square D = b^2 - 120 D = 20^2 - 120 ... plug in b = 20 D = 400 - 120 D = 280 280 is NOT a perfect square, so 10y^2 + 20y + 3 is NOT factorable

jimthompson5910 (jim_thompson5910):

Do the same for the other answer choices until you get a perfect square for D

OpenStudy (anonymous):

Which binomial is a factor of 6y^2 - 13y + 6????

jimthompson5910 (jim_thompson5910):

6y^2-13y+6 6y^2-4y-9y+6 ... Note: -4 and -9 add to -13 and multiply to 6*6 = 36 (6y^2-4y)+(-9y+6) 2y(3y-2)+(-9y+6) 2y(3y-2)-3(3y-2) (2y-3)(3y-2)

jimthompson5910 (jim_thompson5910):

So 6y^2-13y+6 factors to (2y-3)(3y-2)

OpenStudy (anonymous):

I SEEE!!! wow thanks

jimthompson5910 (jim_thompson5910):

yw

OpenStudy (anonymous):

Which binomial is a factor of 24y^2 - 49y + 2?

jimthompson5910 (jim_thompson5910):

24y^2-49y+2 24y^2-y-48y+2 (24y^2-y)+(-48y+2) y(24y-1)+(-48y+2) y(24y-1)-2(24y-1) I'll let you finish up

OpenStudy (anonymous):

wait I don't get where I get y from

jimthompson5910 (jim_thompson5910):

24y^2-y factors to y(24y-1)